Copykwik has a four photocopy machines: A B C and D The probability that a given machine will break down ona particular day is
P(A)= 1/50 P(B)= 1/60 P(C)= 1/75 P(D)= 1/40
assuming independence, what is the probability on a particular da that.
a. All four machines break down?
b. None of the machines will break down?

Answer by stanbon(75887)

Copykwik has a four photocopy machines: A B C and D The probability that a given machine will break down ona particular day is
P(A)= 1/50 P(B)= 1/60 P(C)= 1/75 P(D)= 1/40
assuming independence, what is the probability on a particular da that.
a. All four machines break down?
Multiply all those probabilities to get the answer to "A".
-------------------------------------------------
b. None of the machines will break down?
Subtract the answer of "A" from 1 to get the answer to "B".
===================================================
Cheers,
Stan H.
===========